OpenStudy (nickymail48):

4a^2x^5-6a^5x^2+2a^2x^2 how to factor

OpenStudy (matt101):

\[4a^2x^5-6a^5x^2+2a^2x^2\] You want to look at each factor separately. So starting with the constants, what's the greatest common factor of 4, 6, and 2?

OpenStudy (nickymail48):

2

OpenStudy (funnyguy2):

Let's simplify step-by-step. 4a2x5−6a5x2+2a2x2 =4a2x5+−6a5x2+2a2x2 Answer: =−6a5x2+4a2x5+2a2x2

OpenStudy (matt101):

Right - so we can factor that out of each term: \[2(2a^2x^5-3a^5x^2+a^2x^2)\] Now look at the a's. What's the greatest common factor of a^2, a^5, and a^2?

OpenStudy (nickymail48):

a^2

OpenStudy (matt101):

Right! Now take that out of the 3 terms: \[2a^2(2x^5-3a^3x^2+x^2)\] Now do the same for the x's - greatest common factor of x^5, x^2, and x^2?

OpenStudy (nickymail48):

x^2

OpenStudy (matt101):

Yup...so last step is to take that out to: \[2a^2x^2(2x^3-3a^3+1)\] And you're done!

OpenStudy (nickymail48):

why do the 2x and -3a turn to ^3?

OpenStudy (matt101):

Because: \[a^5 \div a^2=a^3\] Remember your exponent laws: when dividing numbers with the same base (in this case a) but different exponents, you keep the base and subtract the exponents. Same applies for the x's!

OpenStudy (matt101):

When you factor out a number, you divide the terms you're factoring by that number.

OpenStudy (nickymail48):

ok thanks!